Verizon 1:43 AM 47% < Kinematic equations: numerical calcula A basketball rolls onto the court with a speed of 4.0 to the right, and slows down with a constant acceleration of 0.50 over 14 m. What is the velocity of the basketball after rolling for 14m? Round the answer to two significant digits. Answer using a coordinate system where rightward is positive. Stuck? Use a hint. Check agairn
Answer #1

Using third equation of motion,

v^2 = u^2 +2as

v = sqrt(u^2+2as)

= sqrt(4^2 + 2*-0.5*14 )

= 1.4 m/s answer